ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 255, 2023

Agler, J., Ł. Kosiński and J. E. McCarthy Complete normpreserving extensions of holomorphic functions

251

Allen, P., J. Böttcher, D. Clemens and A. Taraz Perfectly packing graphs with bounded degeneracy and many leaves

501

Ansorena, J. L., G. Bello and P. Wojtaszczyk Lorentz spaces and embeddings induced by almost greedy bases in superreflexive Banach spaces

621

Ash, J. M., S. Catoiu and H. Fejzić Gaussian Riemann derivatives  177

Azevedo, J. and P. Shumyatsky Compact groups with probabilistically central monothetic subgroups

955

Bannister, N., J. Bergfalk and J. Tatch Moore On the additivity of strong homology for locally compact separable metric spaces

349

Bao, J. See Z. Liu Bardyla, S. and L. Zdomskyy On regular separable countably

compact ℝ-rigid spaces

783

Bello, G. See J. L. Ansorena Bénard, T. Equidistribution of mass for random processes on finite-

volume spaces

417

Bergfalk, J. See N. Bannister Bispo, S. R. P. and M. Stadlbauer The Martin boundary of an

extension by a hyperbolic group

1

Böttcher, J. See P. Allen Brandenbursky, M. and M. Khanevsky C^{0}-gap between entropy-

zero Hamiltonians and autonomous diffeomorphisms of surfaces

311

Brozos-Vázquez, M., G. Calvaruso, E. García-Río and S. Gavino-Fernández Corrigendum to “Three-dimensional Lorentzian homogeneous Ricci solitons”

975

Bulinski, K. and A. Fish Glasner property for unipotently generated group actions on tori

109

Burness, T. C. and M. Lee On the classification of extremely primitive affine groups

265

Calvaruso, G. See M. Brozos-Vázquez Catoiu, S. See J. M. Ash Chetcuti, E. and B. Zamora-Aviles On the commutant of B(H)

in its ultrapower

423

Clemens, D. See P. Allen Dymond, M. Porosity phenomena of non-expansive, Banach space

mappings

931

García-Río, E. See M. Brozos-Vázquez

Emmanouil, I. K-flatness and orthogonality in homotopy categories  201

Fejzić, H. See J. M. Ash Fish, A. See K. Bulinski Foreman, M., S. Gao, A. Hill, C. E. Silva and B. Weiss Rank-

one transformations, odometers, and finite factors

231

Gallardo-Gutiérrez, E. A., A. G. Siskakis and D. Yakubovich Generators of C_{0}-semigroups of weighted composition operators

63

Gao, S. See M. Foreman Garzoni, D. and N. Gill Large minimal invariable generating sets

in the finite symmetric groups

581

Gavino-Fernández, S. See M. Brozos-Vázquez Gill, N. See D. Garzoni González-Aguilera, V., A. Liendo and P. Montero On the

liftability of the automorphism group of smooth hypersurfaces of the projective space

283

Hill, A. See M. Foreman Juhász, I., J. van Mill, L. Soukup and Z. Szentmiklóssy The

double density spectrum of a topological space

383

Kazhdan, D. and A. Polishchuk Schmidt rank of quartics over perfect fields

851

Khanevsky, M. See M. Brandenbursky Khomskii, Y., M. Koelbing, G. Laguzzi and W. Wohofsky

Laver trees in the generalized Baire space

599

Koelbing, M. See Y. Khomskii Kosiński, Ł. See J. Agler Kucharz, W. and K. Kurdyka A Bochnak–Siciak theorem for Nash

functions over real closed fields

811

Kurdyka, K. See W. Kucharz Laguzzi, G. See Y. Khomskii Lee, M. See T. C. Burness Liendo, A. See V. González-Aguilera

Liu, L. The coding power of a product of partitions  645

Liu, Z. and J. Bao Symmetry of solutions of minimal gradient graph equations on punctured space 81 Liu, Z. S. See Y. Su

Lu, X. and G. Xiong The L_{p} John ellipsoids for negative indices  155
Lytchak, A. and A. Petrunin Weak topology on CAT(0) spaces  763

McCarthy, J. E. See J. Agler van Mill, J. See I. Juhász Miro, E. D., D. Rust, L. Sadun and G. Tadeo Topological mixing

of random substitutions

123

Montero, P. See V. González-Aguilera Narayanan, S., D. Paul and S. Srivastava The multiset partition

algebra

453

Novik, I. and H. Zheng Reconstructing simplicial polytopes from their graphs and affine 2-stresses

891

Papasoglu, P. Polynomial growth and asymptotic dimension  985

Paul, D. See S. Narayanan Petrunin, A. See A. Lytchak Polishchuk, A. See D. Kazhdan

Pournin, L. Shallow sections of the hypercube  685

Rust, D. See E. D. Miro Sadun, L. See E. D. Miro Šaroch, J. Enochs’ conjecture for small precovering classes of mod-

ules

401

Schein, M. M. A family of irreducible supersingular representations of GL_{2}(F) for some ramified p-adic fields

911

Shumyatsky, P. See J. Azevedo

da Silva Jr., G. On the monodromy of elliptic surfaces  871

Silva, C. E. See M. Foreman Siskakis, A. G. See E. A. Gallardo-Gutiérrez Soukup, L. See I. Juhász Srivastava, S. See S. Narayanan Stadlbauer, M. See S. R. P. Bispo Su, Y. and Z. S. Liu Semiclassical states to the nonlinear Choquard

equation with critical growth

729

Szentmiklóssy, Z. See I. Juhász Tadeo, G. See E. D. Miro Taraz, A. See P. Allen Tatch Moore, J. See N. Bannister Weiss, B. See M. Foreman Wohofsky, W. See Y. Khomskii Wojtaszczyk, P. See J. L. Ansorena

Xia, M. On Liu morphisms in non-Archimedean geometry  821

Xiong, G. See X. Lu Yakubovich, D. See E. A. Gallardo-Gutiérrez Yang, D. and Y. Zang Volume growth and topological entropy of

partially hyperbolic systems

325

Yu.Favorov, S. Temperate distributions with locally finite support and spectrum on Euclidean spaces

705

Zamora-Aviles, B. See E. Chetcuti Zang, Y. See D. Yang Zdomskyy, L. See S. Bardyla Zheng, H. See I. Novik